OpenStudy (anonymous):

how do you check if your general solution in DE is correct? I tried solving equations using different methods and plugged in same x and y values. But the arbitrary constant ended up differently.

OpenStudy (akshay_budhkar):

See once you solve a DE you get an arbitrary constant by putting the values of x and y given.. When the values change the value of arbitrary constant also changes.

OpenStudy (akshay_budhkar):

Is that what you asked? or do you have a different query?

OpenStudy (anonymous):

On general principles, if u have found a general solutionto a de, then u should be able to substitute it in the original de to check

OpenStudy (anonymous):

ie not a particular solution

OpenStudy (anonymous):

hmm. but i have rechecked both my solution and nothing seemed wrong. But when I substituted my value of x and y, they both give different constant. Does it mean that my general solution is wrong?

OpenStudy (anonymous):

That's what I am saying, if u substitute the general solution into the de, you can that way guarantee that it right.

OpenStudy (anonymous):

The specifics of a particular solution involve arbitrary constants whose value depends on initial conditions, that is just a calculation.

OpenStudy (anonymous):

Hmm. How do I substitue my general solution in my DE?

OpenStudy (anonymous):

What sort of de is?

OpenStudy (anonymous):

like dy/dx = f(x,y)...

OpenStudy (anonymous):

The only difference between this kind of substitution and the usual sort is that u are substituting a function....

OpenStudy (anonymous):

firstorder >.<

OpenStudy (anonymous):

Do u want an example..?

OpenStudy (anonymous):

Yeah!

OpenStudy (anonymous):

OK, y = 1/2 x^2 + 1/2 and dy/dx = x Simplest one I can think of.....

OpenStudy (anonymous):

SO if u diff the y u get dy/dx = x, right?

OpenStudy (anonymous):

hmmm..

OpenStudy (anonymous):

Of course, your equation probably has more terms....

OpenStudy (anonymous):

How am I going to substitute my general solution to my differential equation? Sorry I still don't get it. >.<

OpenStudy (anonymous):

Should I take the derivative of my gen sol and DE wrt to x/y and see if they are equal?

OpenStudy (anonymous):

Ok, try y = 2e^x -(x^2 +2x +2) and dy/dx = y +x^2

OpenStudy (anonymous):

So diff the y to get dy/dx for one part. Then sub for y in the de and see if both sides are the same.

OpenStudy (anonymous):

If they are, your solution satisfies the de.

OpenStudy (anonymous):

oh so im going to isolate one of my variable and substitue it to DE?

OpenStudy (anonymous):

I'm not sure about the terminology but in the example I just gave u the de is dy/dx = y +x^2 and the solution is y = 2e^x -(x^2 +2x +2) and I want to verify that the solution satisfies the de. So if I differentiate y (my solution) I can get the dy/dx part of my de. And if I substitute 2e^x -(x^2 +2x +2) for y in the de, then see what I have...

OpenStudy (anonymous):

U end up with 2e^x - 2x -2 = 2e^x -2x -2 True.

OpenStudy (anonymous):

oh! thank you very much! i got it nnow!

OpenStudy (anonymous):

:-)